OpenStudy (anonymous):

would someone help me with: (squared)-128(squared)-16

OpenStudy (anonymous):

\[\sqrt{-128} \sqrt{-16}\] is what I meant

OpenStudy (anonymous):

I started with : \[\sqrt{128} x \sqrt{-1} x \sqrt{16} x \sqrt{-1}\] is that correct? then I got : 11.3137085i x 16i

OpenStudy (mertsj):